OpenStudy (anonymous):

Find the slope of the line through the pair of points. (-1/4,1/3) and (-1/2,0) A.3/4 B.-3/4 C.-4/3 D.4/3

OpenStudy (anonymous):

slope = (y2 - y1) / (x2 - x1) = (0 - 1/3) / ((-1/2) - (-1/4)) = (-1/3) / (-1/4) (both negative signs cancel out) = 4/3

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-\frac{1}{4}\quad ,&\frac{1}{3})\quad &(-\frac{1}{2}\quad ,&0) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):